# Addressing divergent representations from causal interventions on neural networks

### v2

@@ Abstract
A common approach to mechanistic interpretability is to causally manipulate model representations via targeted interventions in order to understand what those representations encode. Here we ask whether such interventions create out-of-distribution (divergent) representations, and whether this raises concerns about how faithful their resulting explanations are to the target model in its natural state. First, we demonstrate empirically that common causal intervention techniques often do shift internal representations away from the natural distribution of the target model. Then, we provide a theoretical analysis of two classes of such divergences: ‘harmless’ divergences that occur in the null-space of the weights and from covariance within behavioral decision boundaries, and ‘pernicious’ divergences that activate hidden network pathways and cause dormant behavioral changes. Finally, in an effort to mitigate the pernicious cases, we modify the Counterfactual Latent (CL) loss from Grant (2025) that regularizes interventions to remain closer to the natural distributions, reducing the likelihood of harmful divergences while preserving the interpretive power of interventions. Together, these results highlight a path towards more reliable interpretability methods.
@@ 1 Introduction
A central goal of mechanistic interpretability is to understand what the internal representations of neural networks (NNs) encode and how this gives rise to their behavior. Perhaps the most powerful approach to pursuing this goal is through causal interventions, where methods such as activation patching and Distributed Alignment Search (DAS) directly manipulate internal representations to test how they affect downstream outputs (geiger2021; geiger2023boundless; wu2024alpacadas; wang2022patching; meng2023activpatching; nanda2022attrpatching; csordás2024). Indeed, even correlational interpretability methods such as Sparse Autoencoders (SAEs) and Principal Component Analysis (PCA) often use causal interventions as a final judge for whether the features they identify are truly meaningful (huang2024ravel; dai2024representational). Causal interventions thus occupy a central place in making functional claims about neural circuitry (pearl2010causalmediation; geiger2024causalabstractiontheoreticalfoundation; geiger2025causalunderpin; lampinen2025representation; braun2025functionalvsrepsim).
@@
Figure 1: Causal interventions can recruit hidden circuits that produce misleadingly confirmatory or dormant behavior. (a) For example, consider original pathways (dashed arrows) for two classes A and B that naturally carry activity to different behavioral outputs $y$. In a hypothetical intervention meant to find path A, patching $h^{1}$ with a divergent representation can activate distinct, hidden pathways (solid arrows) that result in misleadingly confirmatory behavior (orange) and/or undetected behavior (red). (b) Now consider 2D projections of the neural activity of $h^{1}$ for a different network that classifies states into one of 10 classes (denoted by hue). Suppose that natural representations (dark points) lie within well-defined decision boundaries (dashed lines) and covary along causal axes, and that intervened representations (light points) are constructed by patching the first axis from a sampled natural representation. Although these representations diverge from the native distribution, this can be harmless (top) or pernicious (bottom) depending on the network’s functional landscape. In particular, it can be pernicious if the network has a functional landscape where intervened activity may unknowingly recruit hidden circuits (visualized as an orange region) or cross dormant behavioral boundaries (red regions).
@@
The use of causal interventions for mechanistic interpretability rests upon a fundamental assumption that counterfactual model states created by interventions are realistic for the target model. Despite its pervasiveness, however, this assumption is often untested. For example, some activation patching experiments multiply feature values by up to 15x (lindsey2025biology); in these settings, it seems possible that intervened representations diverge significantly from the NN’s natural distributions. This raises questions about the reliability of causal interventions for mechanistic interpretability. Do divergent representations change what an intervention can say about an NN’s original mechanisms? When, and to what extent, is it okay for such divergences to occur? When it is not okay, how can we prevent them from occurring?
@@
In this work, we provide both empirical and theoretical insight on these issues. We first demonstrate that divergent representations are a common issue for causal interventions – across a wide range of intervention methods, we find that intervened representations often do diverge from the target NN’s natural distribution. We then provide theoretical examples of two types of divergent representations: ‘harmless’ divergences that can occur from within-decision-boundary covariance along causal dimensions or from deviations along the null-space of the NN layers, which can even be desirable for some mechanistic claims, and ‘pernicious’ divergences that activate hidden network pathways and can cause dormant changes to behavior. We provide discussion on how the harmless and pernicious cases are not always mutually exclusive, where the harm often depends on the specific mechanistic claims. Finally, we provide a broad-stroke, initial solution for mitigating pernicious divergences by minimizing all divergences along causal subspaces (as discovered by DAS). This solution consists of a modified version of the Counterfactual Latent (CL) loss introduced in grant2025mas applied to constrain intervened representations along causal dimensions. We confirm its promise by showing that it can improve out-of-distribution (OOD) performance on interventions in synthetic tasks. Although we do not propose this method as the final solution to representational divergence, we pose it as a step towards more reliable interventions. Our contributions are as follows.
@@
    We show empirical examples of divergence between natural and causally intervened representations for a variety of causal methods (Section 3).
@@
    We provide a theoretical treatment of cases in which divergent representations can arise innocuously from variation along null-space directions and from covariance within decision boundaries of behavioral subspaces, demonstrating that some types of divergence can be okay and even desired (Section 4.1)
@@
    We provide synthetic examples of cases where divergent representations can (1) activate hidden, non-native computational pathways while still resulting in hypothesis-affirming behavior, and (2) cause undetected, latent behavioral changes, together raising questions about the mechanistic claims that can be made from patching results alone (Section 4.2).
@@
    We adapt the CL loss from grant2025mas by averaging over vector labels and exclusively minimizing divergence along causal dimensions. We then use this adapted CL loss to improve OOD performance on causal interventions, providing an initial path towards reducing pernicious divergences (Section 5).
@@ 2.2 Distributed Alignment Search
Distributed Alignment Search (DAS) (geiger2021; geiger2023boundless; wu2024alpacadas) can be understood as a form of activation patching that operates in a transformed basis so that specific, causally relevant subspaces can be manipulated analogously to high-level variables from causal abstractions (CAs) (e.g. symbolic programs). Many cases of individual neuron patching can be understood as a specific case of DAS that uses the identity transform. We use DAS throughout Sections 4 and 5, so here we introduce its theory and background.
DAS is a general framework for finding alignments between subspaces within an NN’s latent vectors and variables from causal abstractions (CAs). DAS achieves this by testing the hypothesis that an NN’s latent state vector $h\in R^{d_{m}}$ can be transformed into a vector $z\in R^{d_{m}}$ that consists of orthogonal subspaces encoding interpretable variables from CAs. This transformation is performed by a learnable, invertible Alignment Function (AF), $z=\mathcal{A}(h)$. We restrict our considerations to orthogonal AFs of the form $\mathcal{A}(h)=Qh$ where $Q\in R^{d_{m}\times d_{m}}$ is an orthogonal matrix, but AFs can in principle take many forms (grant2025das; sutter2025nonlinrepdilemma). This transformation allows us to formulate $h$ in terms of interpretable variables and to manipulate individual variable values.
For a given CA with variables $\text{var}_{i}\in\{\text{var}_{1},\text{var}_{2},...,\text{var}_{n}\}$, DAS tests the hypothesis that $z$ is composed of subspaces $\vec{z}_{\text{var}_{i}}\in R^{d_{\text{var}_{i}}}$ corresponding to each of the variables from the CA. A causally irrelevant subspace $\vec{z}_{\text{extra}}\in R^{d_{\text{extra}}}$ is also included to encode extraneous, functionally irrelevant activity (i.e., the behavioral null-space).
$\mathcal{A}(h)=z=\begin{bmatrix}\vec{z}_{\text{var}_{1}}\\ \vec{z}_{\text{var}_{2}}\\ \cdots\\ \vec{z}_{\text{var}_{n}}\\ \vec{z}_{\text{extra}}\end{bmatrix}$ (1)
@@ 3 Are divergent representations a common phenomenon?
Here we empirically demonstrate how common interpretability methods such as mean difference vector patching (feng2024languagemodelsbindentities), Sparse AutoEncoders (SAEs) (bloom2024saelens), and DAS can result in representations that diverge from the original distribution. The possibility of representational divergence has also been indirectly explored through attention patterns in cases of activation patching that use Gaussian noise corruption (zhang2024bestpracticepatching). Figure 2 shows the top two principal components of the original and intervened latent vectors, distinguished by color. These results show that divergence is a common phenomenon in practice and is not specific to any one method. Even simple methods that patch along a single, mean direction are subject to divergence, despite high behavioral accuracy (feng2024languagemodelsbindentities; geiger2023boundless). We quantify this divergence in Figure 2(a) by computing Earth Mover’s Distance (EMD) (villani2009wassersteinEMD) between all components of the original and intervened activity, and then subtracting a baseline EMD of the original activity compared against itself. See Appendix A.1 for details. Note that these panels do not necessarily imply that the respective methods are invalid or that the claims of the papers are incorrect – this depends on the specific causal claims being made, and we next explore ways in which divergence can be pernicious or harmless to these claims.
@@
Figure 2: Representational divergence is a common issue across various forms of patching. (a) Shows the difference in Earth Mover’s Distance (EMD) between natural compared to intervened latent states and natural latent states compared to themselves. (b)-(d) scatter plots show the top two principal components of natural latent states (orange) and corresponding intervened latent states (blue). All are taken from the transformer residual stream at the intervention positions, and PCA is performed over the combined set of native and intervened latent vectors. Specifically, (b) shows the natural and intervened residual stream from a single layer taken from interventions replicating the results of feng2024languagemodelsbindentities who add and subtract intervention vectors computed as the mean difference between two binding states, (c) shows original vectors and the SAE reconstructed vectors for a single transformer layer using SAELens (bloom2024saelens), and (d) shows projected latent states from before and after interchange interventions for Boundless DAS (wu2024alpacadas).
@@ 4.1 Harmless cases
We first explore cases in which we might consider divergent representations to be harmless to many general computational claims. First among these is a set of cases in which divergence is bottlenecked into the null-space of the next computational layer(s). One could even desire to intervene along this subspace as a causal test to demonstrate that such dimensions are indeed in the null-space. These dimensions can be, for example, the null-space of all subsequent weight matrices, or the space of all negative numbers before a ReLU nonlinearity. If one cares only about the model’s output behavior, any divergence that has no effect on the model’s behavior would therefore be innocuous. Indeed, this covers many present day scenarios. However, it is difficult to define the full behavioral null-space in practice, due to the possibility of dormant behavioral changes, which we discuss in Section 4.2.2.
A less obvious case of harmless divergence can occur from covariance within a network’s decision boundaries. See the top panel of Figure 1(b) for a simulated depiction of harmless divergence caused by covariance along causal dimensions. We see diverging intervened representations created by swapping the first causal dimension of sampled natural latents. However, the divergence remains within the desired decision boundaries, resulting in the appropriate counterfactual behavior.
We expand this intuition using a behaviorally binary subspace as an example, where we define a behaviorally binary subspace as one in which causally impacts the outputs of the next processing layer (e.g., classification labels) only through its sign. Formally, let $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$ denote the next processing layer where $k$ is the output dimensionality. Let $\operatorname{sign}(\cdot)$ denote the elementwise sign map $\operatorname{sign}:\mathbb{R}^{d_{\text{var}}}\to\{-1,0,1\}^{d_{\text{var}}}$, and assume a fixed alignment function $\mathcal{A}$ and subspace selection matrix $D_{\text{var}}\in\mathbb{R}^{d\times d}$. A linear subspace $Z\subseteq\mathbb{R}^{d}$ is behaviorally binary (with respect to $f$, $D_{\text{var}}$, and $\mathcal{A}$) iff for all $D_{\text{var}}\mathcal{A}(h),D_{\text{var}}\mathcal{A}(h^{\prime})\in Z$,
@@
If we focus only on the divergence in causal subspaces and ignore cases of null-space interactions, what claims can we make about the model’s natural neural mechanisms using divergent, intervened representations? An interpretation consistent with the principles of superposition (Smolensky1988; PDP1; PDP2; elhage2022superposition) is that differences in the exact values of each causal subspace do not matter – rather, only the decision boundaries along these subspaces matter. Under this interpretation, any divergence in the intervened distribution arising from covariance in causal subspaces is okay, because the separation of the causal subspaces and their decision boundaries—the computationally important aspects of the NN—are respected in the alignment. This interpretation allows us to ignore computationally irrelevant covariance in our attempt to understand NN mechanisms.

### v3

@@ Abstract
A common approach to mechanistic interpretability is to causally manipulate model representations via targeted interventions in order to understand what those representations encode. Here we ask whether such interventions create out-of-distribution (divergent) representations, and whether this raises concerns about how faithful their resulting explanations are to the target model in its natural state. First, we demonstrate theoretically and empirically that common causal intervention techniques often do shift internal representations away from the natural distribution of the target model. Then, we provide a theoretical analysis of two classes of such divergences: ‘harmless’ divergences that occur in the null-space of the weights and from covariance within behavioral decision boundaries, and ‘pernicious’ divergences that activate hidden network pathways and cause dormant behavioral changes. Finally, in an effort to mitigate the pernicious cases, we apply and modify the Counterfactual Latent (CL) loss from Grant (2025) allowing representations from causal interventions to remain closer to the natural distribution, reducing the likelihood of harmful divergences while preserving the interpretive power of the interventions. Together, these results highlight a path towards more reliable interpretability methods.
 [^1] Link to public repository: https://github.com/grantsrb/rep_divergence
@@ 1 Introduction
A central goal of mechanistic interpretability is to understand what the internal representations of neural networks (NNs) encode and how this gives rise to their behavior. Perhaps the most powerful approach to pursuing this goal is through causal interventions, where methods such as activation patching and Distributed Alignment Search (DAS) directly manipulate internal representations to test how they affect outputs (geiger2021; geiger2023boundless; wu2024alpacadas; wang2022patching; meng2023activpatching; nanda2022attrpatching; csordás2024). Indeed, even correlational methods such as Sparse Autoencoders (SAEs) and Principal Component Analysis (PCA) often use causal interventions as a final judge for whether the features they identify are truly meaningful (huang2024ravel; dai2024representational). Causal interventions thus occupy a central place in making functional claims about neural circuitry (pearl2010causalmediation; geiger2024causalabstractiontheoreticalfoundation; geiger2025causalunderpin; lampinen2025representation; braun2025functionalvsrepsim).
@@
Figure 1: Causal interventions can recruit hidden circuits that produce misleadingly confirmatory or dormant behavior. (a) Consider natural pathways (dashed arrows) for two classes A and B that carry activity to different behavioral outputs $y$. In a hypothetical intervention meant to find path A, patching $h^{1}$ with a divergent representation can activate distinct, hidden pathways (solid arrows) that result in misleadingly confirmatory behavior (orange) and/or undetected behavior (red). (b) Consider 2D projections of the neural activity of $h^{1}$ for a different network that classifies states into one of 10 classes (denoted by hue). Suppose that natural representations (dark points) lie within well-defined decision boundaries (dashed lines) and covary along causal axes, and that intervened representations (light points) are constructed by patching the first axis from a sampled natural representation. Although these representations diverge from the natural distribution, this can be harmless (top) or pernicious (bottom) depending on the network’s functional landscape. In particular, it can be pernicious if the network has a functional landscape where intervened activity unknowingly recruits hidden circuits (visualized as an orange region) or crosses dormant behavioral boundaries (red regions).
@@
The use of causal interventions rests on a fundamental assumption that counterfactual model states created by interventions are realistic for the target model. Despite its pervasiveness, however, this assumption is often untested. For example, some activation patching experiments multiply feature values by up to 15x (lindsey2025biology); in these settings, it seems possible that intervened representations diverge significantly from the NN’s natural distributions. This raises questions about the reliability of causal interventions for mechanistic interpretability. Do divergent representations change what an intervention can say about an NN’s natural mechanisms? When, and to what extent, is it okay for such divergences to occur? When it is not okay, how can we prevent them from occurring?
@@
In this work, we provide both empirical and theoretical insight on these issues. We first demonstrate that divergent representations are a common issue for causal interventions – across a wide range of intervention methods, we find that intervened representations often do diverge from the target NN’s natural distribution. We then provide theoretical examples of two types of divergence: ‘harmless’ divergences that can occur from within-decision-boundary covariance along causal dimensions or from deviations in the null-spaces of the NN layers, and ‘pernicious’ divergences that activate hidden network pathways and can cause dormant changes to behavior. We provide discussion on how harmless and pernicious cases are not always mutually exclusive, where the harm often depends on the specific mechanistic claims. Finally, we provide a broad-stroke, initial solution for mitigating pernicious divergences by minimizing all intervened divergences. We show that we can use the Counterfactual Latent (CL) auxiliary loss introduced in grant2025mas to reduce all representational divergence in the Boundless DAS setting from wu2024alpacadas while maintaining the same behavioral accuracy; and we introduce a modified version of the CL loss that targets causal subspaces and show that it can improve out-of-distribution (OOD) intervention performance on synthetic tasks. Although we do not propose this method as the final solution to representational divergence, we pose it as a step towards more reliable interventions.
We summarize our contributions as follows:
@@
    We show theoretical and empirical examples of divergence between natural and causally intervened representations for a variety of causal methods (Section 3).
@@
    We provide a theoretical treatment of cases in which divergence can arise innocuously from variation in the null-space and from covariance within decision boundaries of behavioral subspaces, demonstrating that some divergences can be okay and even desired (Section 4.1)
@@
    We provide synthetic examples of cases where divergent representations can (1) activate hidden computational pathways while still resulting in hypothesis-affirming behavior, and (2) cause dormant behavioral changes, together raising questions about the mechanistic claims that can be made from patching results alone (Section 4.2).
@@
    We use the CL auxiliary loss from grant2025mas to minimize patching divergence directly in the Large Language Model (LLM) Boundless DAS experimental setting from wu2024alpacadas, and we introduce a stand-alone, modified CL loss that exclusively minimizes divergence along causal dimensions, improving OOD intervention performance in synthetic settings. Together, these results provide an initial step towards preventing pernicious divergences (Section 5).
@@ 2.2 Distributed Alignment Search
Distributed Alignment Search (DAS) (geiger2021; geiger2023boundless; wu2024alpacadas) can be understood as a form of activation patching that operates in a transformed basis so that specific, causally relevant subspaces can be manipulated analogously to high-level variables from causal abstractions (CAs) (e.g. symbolic programs). Many cases of individual neuron (coordinate) patching can be understood as specific cases of DAS that use the identity transform. We use DAS in Sections 4 and 5, so here we introduce its theory and background.
DAS finds alignments between neural representational subspaces and causal variables from Causal Abstractions (CAs) by testing the hypothesis that an NN’s latent state vector $h\in R^{d_{m}}$ can be transformed into a vector $z\in R^{d_{m}}$ that consists of orthogonal subspaces encoding interpretable variables. This transformation is performed by a learnable, invertible Alignment Function (AF), $z=\mathcal{A}(h)$. We restrict our considerations to linear AFs of the form $\mathcal{A}(h)=Wh$ where $W\in R^{d_{m}\times d_{m}}$ is invertible. This transformation allows us to formulate $h$ in terms of interpretable variables and to manipulate encoded values.
For a given CA with variables $\text{var}_{i}\in\{\text{var}_{1},\text{var}_{2},...,\text{var}_{n}\}$, DAS tests the hypothesis that $z$ is composed of subspaces $\vec{z}_{\text{var}_{i}}\in R^{d_{\text{var}_{i}}}$ corresponding to each of the variables from the CA. A causally irrelevant subspace $\vec{z}_{\text{extra}}\in R^{d_{\text{extra}}}$ is also included to encode extraneous, functionally irrelevant neural activity (i.e., the behavioral null-space).
$\mathcal{A}(h)=z=\begin{bmatrix}\vec{z}_{\text{var}_{1}}\\ \cdots\\ \vec{z}_{\text{var}_{n}}\\ \vec{z}_{\text{extra}}\end{bmatrix}$ (1)
@@ 3 Are divergent representations a common phenomenon?
We begin by demonstrating that divergent representations are a common (if not likely) outcome of causal interventions, both in theory and in practice. We do not yet consider its perniciousness, however, and we reserve the question of whether and when divergence is harmful for Section 4.
3.1 For most manifolds, coordinate patching guarantees divergence
We first consider a theoretical setting where coordinate-based patching of one or more vector dimensions is performed on a single manifold, similar to what might be done in neuron level activation patching (vig2020causal; geiger2021). We prove that in this setting, divergent representations are guaranteed to occur if patching is performed exhaustively. For simplicity, we consider a minimal version of this proof that involves a circular manifold with two dimensions. A more general proof, which applies to most manifold geometries, can be found in Appendix A.2.
@@ 3.1 For most manifolds, coordinate patching guarantees divergence
Figure 2: Representational divergence is a common occurrence across various interventions. (a) Directly replacing a coordinate value in one natural representation (orange) with the value from another will eventually create divergent representations (blue). (b) Top two principal components of natural and corresponding intervened representations, taken from the residual stream at the intervention position and with PCA is performed over the combined set of natural and intervened vectors, for three popular causal intervention techniques: a replication of feng2024languagemodelsbindentities for mean difference patching, reconstructed vectors for a single transformer layer using SAELens (bloom2024saelens) for sparse autoencoder, and interchange interventions for Boundless DAS (wu2024alpacadas). (c) L2 distance between natural and corresponding intervened representations, and Earth Mover’s Distance (EMD) between natural and intervened distributions (with baseline comparing the natural distribution to itself).
Formally, let $\mathcal{M}_{K}=\{\,c_{K}+u:\|u\|_{2}\leq r_{K}\,\}\subset\mathbb{R}^{2}$ be a class-$K$ manifold with centroid $c_{K}\in\mathbb{R}^{2}$ and radius $r_{K}>0$. Given two native representations $h^{\textit{src}}=c_{K}+u$ and $h^{\textit{trg}}=c_{K}+v$, let us define a coordinate patch (onto class $K$) that keeps the first coordinate of $h^{\textit{src}}$ and the second of $h^{\textit{trg}}$:
$\hat{h}\;=\;\begin{bmatrix}h_{1}^{\textit{src}}\\ h_{2}^{\textit{trg}}\end{bmatrix}=\begin{bmatrix}c_{K,1}+u_{1}\\ c_{K,2}+v_{2}\end{bmatrix}$.
Then the deviation from $c_{K}$ is
$\hat{h}-c_{K}\;=\;(u_{1},\,v_{2})^{\top},\qquad\|\hat{h}-c_{K}\|_{2}^{2}\;=\;u_{1}^{2}+v_{2}^{2}.$
Proposition (coordinate patching exceeds the class radius). If $h^{\textit{src}},h^{\textit{trg}}\in\mathcal{M}_{K}$ (i.e., $\|u\|_{2}\leq r_{K}$ and $\|v\|_{2}\leq r_{K}$), then the patched point $\hat{h}$ is off-manifold whenever $u_{1}^{2}+v_{2}^{2}>r_{K}^{2}$. In particular, there exist boundary points $h^{\textit{src}},h^{\textit{trg}}\in\partial\mathcal{M}_{K}$ with $u=(r_{K},0)$ and $v=(0,r_{K})$ such that $\|\hat{h}-c_{K}\|_{2}=\sqrt{r_{K}^{2}+r_{K}^{2}}=r_{K}\sqrt{2}>r_{K}$.
Proof. Since $\hat{h}-c_{K}=(u_{1},v_{2})$, we have $\|\hat{h}-c_{K}\|_{2}^{2}=u_{1}^{2}+v_{2}^{2}$. Choosing $u=(r_{K},0)$ and $v=(0,r_{K})$ gives the stated violation. ∎
As noted above, this intuition holds for all manifold shapes other than axis-aligned hyper-rectangles (see Appendix A.2). Thus, in these relatively simple theoretical intervention settings, divergent representations are guaranteed to occur with enough intervention samples.
3.2 Many existing causal methods empirically produce divergence
Stepping beyond the theoretical setting, we also empirically demonstrate that common, real-world causal interpretability methods often produce divergent representations. The notion that these methods can create divergence has been indirectly explored in previous work through attention patterns in Gaussian noise corrupted activation patching (zhang2024bestpracticepatching). Here we extend this view to three popular causal intervention methods: mean difference vector patching (e.g. feng2024languagemodelsbindentities), Sparse Autoencoders (e.g. bloom2024saelens), and DAS (e.g. wu2024alpacadas).
Figure 2 shows the top two principal components of the natural and intervened representations for each intervention method, distinguished by color. These results demonstrate that divergence is a common phenomenon in practice and is not specific to any one method. Even simple methods that patch along a single mean direction are subject to divergence, despite high behavioral accuracy (feng2024languagemodelsbindentities; geiger2023boundless; wu2024alpacadas). We quantify this divergence in Figure 2(c) using the Earth Mover’s Distance (EMD) (villani2009wassersteinEMD) between the full dimensional natural and intervened distributions, using the corresponding natural-natural comparison as a baseline. We see that the intervened divergence exceeds that of the natural (more metrics and details are in Appendix A.1). Note that this result does not necessarily imply that the respective methods are invalid or that their claims are incorrect; the panels are only meant to show the presence of divergence.
@@ 4 When are divergent representations harmless or pernicious?
Having demonstrated that divergent representations are a common phenomenon, we now turn our attention to considering whether divergence is a concerning phenomenon. We propose that divergence is harmless to most functional claims if it occurs in the behavioral null-space of the model weights or from covariance within decision boundaries, and that it can be pernicious if it recruits hidden pathways or leads to dormant behavioral changes. However, we stress that the harm is inherently claim dependent, meaning that these divergences are not mutually exclusive.
@@ 4.1 Harmless cases
We first explore cases in which we might consider divergent representations to be harmless to many general functional claims. First among these is a set of cases in which divergence is bottlenecked into the null-space of the next computational layer(s). Formally, we define the null-space of a weight matrix $W\in\mathbb{R}^{d_{2}\times d_{1}}$ as $\mathcal{N}(W)=\{v\in\mathbb{R}^{d_{1}}\,\,|\,\,Wv=0\}$. Neural activity in the null-space of the weights refers to any changes $\delta\in\mathbb{R}^{d_{1}}$ for which $W(h^{\ell}+\delta)=Wh^{\ell}$. We can generalize this notion beyond the weights to all remaining computations after and including layer $\ell$ in a neural network denoted $f_{\geq\ell}(h)$. We call this the behavioral null-space for a layer $\ell$, defined as $\mathcal{N}(f_{\geq\ell}(h^{\ell}))=\{v\in\mathbb{R}^{d_{1}}\,\,|\,\,f_{\geq\ell}(h^{\ell}+v)=f_{\geq\ell}(h^{\ell})\}$.
If one cares only about the model’s output behavior, then divergence that is bottlenecked into null-spaces and that therefore has no effect on the model’s behavior would be harmless. Indeed, this covers many present day scenarios, and one could even desire to intervene along null-space directions in order to causally test that such dimensions are indeed in the null-space. It is, however, difficult to define the full behavioral null-space in practice due to the possibility of dormant behavioral changes, which we discuss in Section 4.2.2.
A less obvious case of harmless divergence can occur from covariance within a network’s decision boundaries. For example, the top panel of Figure 1(b) provides a simulated depiction of harmless divergence caused by covariance along causal dimensions. Here, divergent representations are created by swapping the first causal dimension of sampled natural vectors, however, these representations remain within the desired decision boundaries, resulting in the appropriate counterfactual behavior.
We can expand this intuition using a behaviorally binary subspace as an example, where we define a behaviorally binary subspace as a vector subspace which causally impacts the outputs of a future processing layer (e.g., classification labels) only through its sign. Formally, let $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$ denote a computational layer (possibly consisting of multiple NN layers and functions) where $k$ is the output dimensionality. Let $\operatorname{sign}(\cdot)$ denote the elementwise sign map $\operatorname{sign}:\mathbb{R}^{d_{\text{var}}}\to\{-1,1\}^{d_{\text{var}}}$, and assume a fixed alignment function $\mathcal{A}$ and subspace selection matrix $D_{\text{var}}\in\mathbb{R}^{d\times d}$. A linear subspace $Z\subseteq\mathbb{R}^{d}$ is behaviorally binary (with respect to $f$, $D_{\text{var}}$, and $\mathcal{A}$) iff for all $D_{\text{var}}\mathcal{A}(h),D_{\text{var}}\mathcal{A}(h^{\prime})\in Z$,
@@
If we focus only on the divergence in causal subspaces and ignore cases of null-space interactions, what claims can we make about the model’s natural neural mechanisms using divergent, intervened representations? An interpretation consistent with the principles of superposition (Smolensky1988; PDP1; PDP2; elhage2022superposition) is that differences in the exact values of each causal subspace do not matter – rather, only the decision boundaries along these subspaces matter. Under this interpretation, any divergence in the intervened distribution arising from covariance in causal subspaces is acceptable because the separation of the causal subspaces and their decision boundaries (i.e. the computationally important aspects of the NN) are respected in the alignment. This interpretation allows us to ignore computationally irrelevant covariance in our attempt to understand NN mechanisms.